a) The Ghana Economic Management Association wishes to have information on the mean income of store managers in the retail industry. A random sample of 256 managers reveals a sample mean of Ghc45,420. The standard deviation of this population is Ghc2,050. What is a reasonable range of values for the population mean? Interpret your results. 

b) i. Briefly outline the steps involved in hypothesis testing 

ii. Use a table to illustrate how Type I and Type II error can occur


1
Expert's answer
2021-08-13T08:50:38-0400

Part a

90% CI for "\mu" using Normal dist

Sample Mean "| \bar{x} = 45420"

Since the sample size is large, sample standard deviation can be used to approximate population standard deviation.

Population Standard deviation "=|\sigma =2050"

Sample Size "=| n = 256"

Significance level "= \alpha =1— Confidence =1-0.9 = 0.1"

The Critical Value"= z_{\alpha/2} =z_{0.06} =1.645" (From z table , using interpolation "\frac{1}{2}" the distance between 164 and 1.65 )

Critical Values , "± z_{\alpha/ 2}=±1.645"

Margin of Error "= E = z_{\alpha/2}*\frac{\sigma}{\sqrt{n}} =1.645 *\frac{2050}{\sqrt{256}}=210.765625"

Margin of Error, "E= 210.766"

Limits of 90% confidence interval are given by

Lower limit "= \bar{x} - E = 45420 — 210.365625 = 45209.234375 = 45 209.234"

Upper limit "=\bar{x}+ E = 45420 +210.765625 = 45630.365625 = 45 630.766"  

90% confidence interval is "\bar{x}±E = 45420± 210.765625 = (45209.234375, 45630.765625)"

90% CI using normal dist: "(45 202.234) < \mu < (45 630.766)"


95% CI for "\mu" using Normal dist

Sample Mean "| \bar{x} = 45420"

Since the sample size is large, sample standard deviation can be used to approximate population standard deviation.

Population Standard deviation "=|\sigma =2050"

Sample Size "=| n = 256"

Significance level "= \alpha =1— Confidence =1-0.95 = 0.05"

The Critical Value"= z_{\alpha/2} =z_{0.025} =1.96"

Critical Values , "± z_{\alpha/ 2}=±1.96"

Margin of Error "= E = z_{\alpha/2}*\frac{\sigma}{\sqrt{n}} =1.96 *\frac{2050}{\sqrt{256}}=251.125"

Margin of Error, "E= 251.125"

Limits of 95% confidence interval are given by

Lower limit "= \bar{x} - E = 45420 - 251.125 = 45168.875"

Upper limit "=\bar{x}+ E = 45420 +251.125 = 45671.125"  

95% confidence interval is "\bar{x}±E = 45420± 251.125 = (45168.875,45671.125)"

90% CI using normal dist: "(45168.875) < \mu < (45671.125)"


99% CI for "\mu" using Normal dist

Sample Mean "| \bar{x} = 45420"

Population Standard deviation "=|\sigma =2050"

Sample Size "=| n = 256"

Significance level "= \alpha =1— Confidence =1-0.99 = 0.01"

The Critical Value"= z_{\alpha/2} =z_{0.005} =2.58"

Critical Values , "± z_{\alpha/ 2}=±2.58"

Margin of Error "= E = z_{\alpha/2}*\frac{\sigma}{\sqrt{n}} =2.58 *\frac{2050}{\sqrt{256}}=330.5625"

Margin of Error, "E= 330.5625"

Limits of 99% confidence interval are given by

Lower limit "= \bar{x} - E = 45420 - 303.563 = 45089.438"

Upper limit "=\bar{x}+ E = 45420 +303.5625= 45750.563"  

99% confidence interval is "\bar{x}±E = 45420± 303.5625 = (45089.438,45750.503)"

99% CI using normal dist: "(45089.438) < \mu < (45750.503)"

The widest CI is the one with highest confidence level. 

99% CI is the widest. 


Part b

i)

Enter the Null Hypothesis.

Describe the Alternative Hypothesis.

Determine the Significance Level (a)

Determine the Test Statistic and the Corresponding P-Value.

Making a decision.

ii)

In statistics, a Type I error means rejecting the null hypothesis when it's actually true. In contrast, a Type II error means failing to reject the null hypothesis when it's actually false.